OpenStudy (nikkiv73):

Find the average rate of change of the function defined by the following table. x y -2 -5 2 35 6 75 10 115 -10, 1/10, -1/10, 10 - 10

OpenStudy (.mehek.):

Do you have any idea what the answer might be?

OpenStudy (nikkiv73):

10

OpenStudy (nikkiv73):

slope = rise/run

OpenStudy (nikkiv73):

but i still dont get it

OpenStudy (.mehek.):

Correct

OpenStudy (.mehek.):

You can pick 2 points and use the slope formula

OpenStudy (sadlalkdja12111312):

y increase by 40 and x increase by 4

OpenStudy (.mehek.):

\(\dfrac{35-(-5)}{2-(-2)}=\dfrac{40}{4}=10\) @nikkiv73 was correct

OpenStudy (nikkiv73):

Thank you! I need more coffee.